# Data-driven model order reduction for wave propagation in materials with nonlinearities or damage

### v1

@@ 4.2 Numerical results for the mechanical system resulting from GUW propagation
In this subsection, we present the numerical results for the application of the non-intrusive model order reduction techniques operator inference (3.2, 3.4, 3.5) and dynamic mode decomposition (3.6) to the problem of the mechanical system (11) obtained upon the discretization of the equation of motion (2) for elasticity problems. We consider the two-dimensional physical model of a fiber metal laminate (FML) of [BellamMuralidhar2021]. The model contains one physical defect and describes the propagation of an anti-symmetric lamb wave mode ($A_{0}$) in the FML and the subsequent interaction with the defect. A sketch of the model is shown in Figure 7. The FML comprises sixteen alternating layers of carbon fiber reinforced plastic (CFRP) and steel. Damage is introduced by locally reducing the Young’s modulus $E_{d}$ in one of the steel layers. The simulation of the GUW propagation and its interaction with the damage is carried out using the commercial software COMSOL-Multiphysics software®. The model is excited on the top and bottom left nodes of the model with a force excitation using a five-cycle Hanning window sinusoidal burst with a central frequency of $120$ kHz. The model is excited for $4.167\times 10^{-5}$ $\mathrm{s}$ and the total time of the simulation is $2.083\times 10^{-4}$ $\mathrm{s}$. The thicknesses of steel and CFRP layers are $0.12$ $\mathrm{m}$ and $0.125$ $\mathrm{m}$, respectively. The spatial discretization yields $N=79266$ degrees of freedom. The location of the damage along the FML is called $x_{d}$ and the length of the damage along the FML is denoted by $l_{d}$. The excitation force as function of time is shown in Figure 8. The time step size of the numerical simulation is $2.0833\times 10^{-7}$, the solution is gathered at each time step leading to $M=1001$ snapshots. We present results for the damage configuration with the Young’s modulus of elasticity $E_{d}=2.55\times 10^{9}$ $\mathrm{P}\mathrm{a}$, the position of the damage $x_{d}=0.06$ $\mathrm{m}$ and the length of the damage $l_{d}=0.002$ $\mathrm{m}$. A sensor is placed in the domain measuring the out-of-plane displacement (vertical displacement).
@@
In this numerical example, we compare the DMD method and its multiresolution variant (mrDMD) to the operator inference approach. For DMD, the implementation details are given as follows. Like in the first example, we use two different strategies to compute the initial mode amplitudes: (i) projection of the first snapshot onto the POD basis (see (46)), and (ii) an optimized amplitude formulation based on a least-squares fit over the time series (see (47)). In addition to the classical DMD, we use mrDMD which aggregates modal contributions from successive time-bins at different scales, ultimately producing a multiscale reconstruction of the original dynamics. For mrDMD, the maximum number of hierarchical time levels is set to $6$. This decomposition technique requires specifying a frequency threshold to distinguish between slow and fast dynamical modes. In this example, no slow modes are identified at the first level, which spans the full time window. At the second level, six significant modes are extracted. The temporal evolution of their coefficients, defined by $\mathbf{\Psi}=\mathbf{b}\,\exp(\Omega t)$, is illustrated in Figure 9 for the second level dynamics which contains $4$ different time-bins, the number of extracted modes here is $6$ corresponding to $3$ different pairs of frequencies.
@@
Figure 10 (left) compares the full order model (FOM) and DMD predictions at the sensor location. The DMD signal is obtained using (41), whereas the DMD Optimal one is computed using (49) and the reduced dimension is set to $80$. The DMD reconstruction using the first snapshot for amplitude estimation performed poorly, failing to recover the dynamics. This is attributed to the fact that the initial displacement field is nearly zero everywhere except at the excitation points, leading to negligible initial amplitudes. In contrast, DMD with optimized amplitudes accurately captured the signal, yielding a relative error (see (51)) at the sensor of $\epsilon_{s}=0.626\%$.
@@
The mrDMD reconstruction closely matched the FOM signal as well, as shown in Figure 10 (right), achieving a lower sensor error of $\epsilon_{s}=0.137\%$. This demonstrates that the multiresolution approach is effective at isolating time-localized dynamics in this guided wave simulation.
@@
Finally, we compare the computational costs. DMD with optimized amplitudes and rank $r=80$ required $5.511$ $\mathrm{s}$ for the computation of the DMD modes and eigenvalues, the construction of the Vandermode matrix and the resolution of the linear system for the optimal amplitudes vector $\mathbf{b}$. Then only $0.456$ $\mathrm{s}$ are needed to compute the time dynamics $\mathbf{\Psi}$ and the reconstruction of the reduced solution. In the case of mrDMD, the method required $8.919$ $\mathrm{s}$ for the computation of the DMD modes over the different time levels and time-bins inside each level, additional $3.218$ $\mathrm{s}$ were taken for the extraction of the time dynamics and DMD modes from each time-bin and finally reconstructing the mrDMD solution. Although mrDMD offered improved accuracy, the classical DMD was computationally more efficient.
Refer to captionRefer to caption
@@
Figure 10: Comparison between the FOM and DMD predicted displacement (left) and the mrDMD predicted displacement (right) at the sensor for the mechanical model of the damaged FML. The reduced dimension is $r=80$ in the case of DMD and DMD Optimal.
@@
Next, we apply the operator inference approach to this model. We recall that in order to apply operator inference one has to obtain the approximation of the second derivative data. In our case, the COMSOL software provides only the data for the solution and its first time derivative. Therefore, we employ an eighth order finite difference scheme, see [Sharma2024], in order to approximate accurately the second time derivatives of the solution. Then we proceed with obtaining the reduced basis by carrying out a SVD of the snapshot matrix and thereafter we constructed reduced versions of the displacement and its derivatives. Figure 11 (left) shows the decay of the singular values of the snapshot matrix of the displacement. The cumulative sum of the singular values is shown in Figure 11 (right). We observe that at least the first $50$ singular values are needed for recovering $98$ $\%$ of the energy embedded in the snapshot data.
Refer to captionRefer to caption
@@
The operator inference approach used for this example is the unconstrained one which does not assume full knowledge about the external loads, see Section 3.4. In this approach, we assume only the presence of the input signal (the excitation signal) mentioned earlier. Therefore, the dimension of the control matrix $\widehat{\mathbf{B}}_{M}$ on the right hand side of the reduced system is $r\times 1$.
@@
In this example, we have different scales of the snapshot matrices and the input signal. Therefore, we apply the scaling pre-processing step as explained in Section 3.5. The solution of the optimization problem of the OpInf was obtained using the stochastic gradient optimizer ADAM [Adam] from the software library of TensorFlow [TensorFlow], a Tikhonov regularization term proportional to the Frobenius norm of the inferred operators is added to stabilize the solution. We infer the reduced operators for different sizes of the reduced model in order to study the effect of the reduced dimension on the quality of the approximation. In particular, we simulate the OpInf problem for $r=50,60,70,80,100$ and $150$. We obtained the reduced operators of $\widehat{\mathbf{K}}_{M}$ and $\widehat{\mathbf{B}}_{M}$ for each setting. These reduced operators are then used to integrate the second order mechanical system using the Newmark method. The solution obtained by the OpInf based ROM and the reference solution of COMSOL are plotted at the sensor location in Figure 12. The plots indicate that the OpInf approach allows for qualitatively accurate approximations of the reference solution for $r=50$ and it can be seen that the reduced approximation is improved when considering $r=80$. A more quantitative assessment of the approach is provided by computing the relative error at the sensor $\epsilon_{s}$ as defined in (51). To analyze the convergence of the OpInf reduced approximation with respect to the reduced dimension, we compute the values of $\epsilon_{s}$ for increasing numbers of reduced basis functions. The values of the sensor error are plotted versus the reduced dimension $r$ in Figure 13 (left) showing that the error reaches smaller values when increasing the reduced dimension $r$.
@@
To further evaluate the quality of the reduced approximation, we also examine the time-dependent relative error over the spatial domain $\epsilon_{u}(t)$, defined in (52). The error $\epsilon_{u}(t)$ for $r=80$ is presented in Figure 13 (right). As it can be seen from the plot, the spatial error $\epsilon_{u}(t)$ remains under the threshold of 55 $\%$ for the most part of the simulation which shows that the OpInf-ROM is able to contain the error growth.
@@
Figure 12: The reduced order model solution obtained via unconstrained OpInf for $r=50$ (left) and $r=80$ (right) versus the reference solution of COMSOL at the sensor for the mechanical model of the damaged FML.
@@ 4.3 Numerical results for the nonlinear hyperelastic model
The third numerical example investigates a nonlinear elastic model of an aluminium plate. The model consists of a two-dimensional, intact rectangular aluminium plate with dimensions $0.3$ $\mathrm{m}$ (length) × $0.004$ $\mathrm{m}$ (thickness). A sketch of the model is shown in Figure 14. We note that this model has been investigated in our article [GH24], where we considered antisymmetric lamb wave A0. In contrast, here we consider a symmetric lamb wave mode S0. Actuators are positioned at the top and bottom left corners of the plate, applying a five-cycle Hanning window sinusoidal burst with a central frequency of $100$ kHz. The material is modeled as hyperelastic, with nonlinearity captured using the Neo-Hookean model [Rivlin1948]. The model employs displacement-based excitation, implemented as nonhomogeneous Dirichlet boundary conditions for the vertical displacement at the two left ends of the plate. A homogeneous Dirichlet boundary condition is set for the mid-point on the right side of the plate. The simulation is run for $2.5\times 10^{-4}$ $\mathrm{s}$, with excitation active during the initial $5\times 10^{-5}$ $\mathrm{s}$. The simulation of the GUW is obtained by COMSOL-Multiphysics software®.
Figure 14: The aluminium model setup for the GUW simulation. The red arrows are used to indicate the sign of the excitation signal at each of the two nodes. A symmetric excitation is used for this numerical example through the displacement.
@@
Standard DMD with a reduced dimension of $r=80$ is applied to the same snapshot set, with initial amplitudes computed using Equations (46) and (47). However, the method fails in both cases, as demonstrated in Figure 17. This figure compares the out-of-plane displacement measured by the sensor for the reference COMSOL solution and the DMD approximations; standard DMD does not provide even qualitatively accurate results. This does not change as we increase the reduced dimension $r$. In contrast, Figure 17 shows a similar comparison using mrDMD, where the alignment between the signals is significantly better. The relative error $\epsilon_{s}$ for the mrDMD approximation is $1.023$ $\%$, which is within an acceptable range.
@@ 5 Outlook
In this work, we apply both intrusive and non-intrusive model reduction approaches to the wave equation posed on a domain containing damage. Subsequently, we apply non-intrusive reduction techniques to discretized second order mechanical systems obtained from guided ultrasonic wave (GUW) simulations. The test cases considered include models exhibiting either damage or nonlinear behavior.
Damage identification and characterization—determining both the location and severity of defects—are central challenges in the structural health monitoring (SHM) community. This motivation underlies our current research, which aims to develop a reduced order model (ROM) capable of generalizing across different damage configurations. To achieve this, we have to parameterize the damage and address the associated inverse problem, which involves estimating unknown damage parameters from observed system outputs.
As discussed in this article, the input damage parameters influence the full order stiffness matrix operator and, consequently, the displacement field. This dependency is carried over to the projected reduced operator of the stiffness matrix. Hence, a key challenge arises in approximating the reduced stiffness operator as a function of the parameters in a fully reduced fashion–i.e., without accessing the full order model or computing any full order quantities during the online stage of the ROM.
In the literature, this challenge is commonly addressed using hyper-reduction techniques such as the Empirical Interpolation Method (EIM) or the Discrete Empirical Interpolation Method (DEIM). For instance, DEIM evaluates a small subset of entries of the full order matrix during the online stage, where the corresponding indices are selected in a greedy manner. However, applying such methods requires that the full order solver or computational code allows direct computation of specific entries of the stiffness matrix without assembling the entire matrix. This is not feasible in our case, since we employ COMSOL Multiphysics as the simulator for the GUW problem, and direct access to the full order matrices or the time-discretization scheme is restricted.
Consequently, we currently adopt a non-intrusive yet projection-based reduction approach–namely, operator inference (OpInf). This choice is motivated by the need to construct a reduced representation of the second order mechanical system governing the dynamics without requiring intrusive modifications to the full order solver. To achieve the objectives of this work, we have to extend the existing OpInf-based ROM framework to handle parameterized systems. The proposed extension must be designed to accurately capture the dependence of the system outputs—in this case, the displacement field—on the damage-related input parameters.

### v2

@@ 4.2 Numerical results for the mechanical system resulting from GUW propagation
In this subsection, we present the numerical results for the application of the non-intrusive model order reduction techniques operator inference (3.3, 3.4, 3.5) and dynamic mode decomposition (3.8) to the problem of the mechanical system (11) obtained upon the discretization of the equation of motion (2) for elasticity problems. We consider the two-dimensional physical model of a fiber metal laminate (FML) of [56]. The model contains one physical defect and describes the propagation of an anti-symmetric lamb wave mode ($A_{0}$) in the FML and the subsequent interaction with the defect. A sketch of the model is shown in Fig. 9. The FML comprises sixteen alternating layers of carbon fiber reinforced plastic (CFRP) and steel. Damage is introduced by locally reducing the Young’s modulus $E_{d}$ in one of the steel layers. The simulation of the GUW propagation and its interaction with the damage is carried out using the commercial software COMSOL-Multiphysics software©. The model is excited on the top and bottom left nodes of the model with a force excitation using a five-cycle Hanning window sinusoidal burst with a central frequency of $120$ kHz. The model is excited for $4.167\times 10^{-5}$ $\mathrm{s}$ and the total time of the simulation is $2.083\times 10^{-4}$ $\mathrm{s}$. The thicknesses of steel and CFRP layers are $0.12$ $\mathrm{m}$ and $0.125$ $\mathrm{m}$, respectively. The spatial discretization yields $N=79266$ degrees of freedom. The location of the damage along the FML is called $x_{d}$ and the length of the damage along the FML is denoted by $l_{d}$. The excitation force as function of time is shown in Fig. 10. The time step size of the numerical simulation is $2.0833\times 10^{-7}$, the solution is gathered at each time step leading to $M=1001$ snapshots. We present results for the damage configuration with the Young’s modulus of elasticity $E_{d}=2.55\times 10^{9}$ $\mathrm{P}\mathrm{a}$, the position of the damage $x_{d}=0.06$ $\mathrm{m}$ and the length of the damage $l_{d}=0.002$ $\mathrm{m}$. A sensor is placed in the domain measuring the out-of-plane displacement (vertical displacement).
@@
The access to the FOM is restricted in this numerical example. Therefore, we compare the DMD method and its multiresolution variant (mrDMD) to the operator inference approach. For DMD, the implementation details are given as follows. Like in the first example, we use two different strategies to compute the initial mode amplitudes: (i) projection of the first snapshot onto the POD basis (see (61)), and (ii) an optimized amplitude formulation based on a least-squares fit over the time series (see (62)). In addition to the classical DMD, we use mrDMD which aggregates modal contributions from successive time-bins at different scales, ultimately producing a multiscale reconstruction of the original dynamics. For mrDMD, the number of hierarchical time levels is set to $6$. This decomposition technique requires specifying a frequency threshold to distinguish between slow and fast dynamical modes. In this example, no slow modes are identified at the first level, which spans the full time window. At the second level, six significant modes are extracted. The temporal evolution of their coefficients, defined by $\mathbf{\Psi}=\mathbf{b}\,\exp(\Omega t)$, is illustrated in Fig. 11 for the second level dynamics which contains $4$ different time-bins, the number of extracted modes here is $6$ corresponding to $3$ different pairs of frequencies.
@@
Figure 12 (left) compares the full order model (FOM) and DMD predictions at the sensor location. The DMD signal is obtained using (56), whereas the optimal DMD is computed using (64) and the reduced dimension is set to $r=80$. The DMD reconstruction using the first snapshot for amplitude estimation performed poorly, failing to recover the dynamics. This is attributed to the fact that the initial displacement field is nearly zero everywhere except at the excitation points, leading to negligible initial amplitudes. In contrast, DMD with optimized amplitudes accurately captured the signal, yielding a relative error (66) at the sensor of $\epsilon_{s}=0.626\%$.
@@
The mrDMD reconstruction closely matched the FOM signal, see Fig. 12 (right), achieving a lower sensor error of $\epsilon_{s}=0.137\%$. This demonstrates that the multiresolution approach is effective at isolating time-localized dynamics in this guided wave simulation.
@@
Finally, we compare the computational costs. DMD with optimized amplitudes and rank $r=80$ required $5.511$ $\mathrm{s}$ for the computation of the DMD modes and eigenvalues, the construction of the Vandermode matrix and the resolution of the linear system for the optimal amplitudes vector $\mathbf{b}$. Then only $0.456$ $\mathrm{s}$ are needed to compute the time dynamics $\mathbf{\Psi}$ and the reconstruction of the reduced solution. In the case of mrDMD, the method required $8.919$ $\mathrm{s}$ for the computation of the DMD modes over the different time levels and time-bins inside each level, additional $3.218$ $\mathrm{s}$ were taken for the extraction of the time dynamics and DMD modes from each time-bin and finally reconstructing the mrDMD solution. Although mrDMD offered improved accuracy, the optimal DMD was computationally more efficient.
@@
Figure 12: Comparison between the FOM and DMD predicted displacement (left) and the mrDMD predicted displacement (right) at the sensor for the mechanical model of the damaged FML. The reduced dimension is $r=80$ in the case of DMD and optimal DMD.
@@
Next, we apply time continuous operator inference approach to this model. We recall that in order to apply operator inference one has to obtain the approximation of the second derivative data. In our case, the COMSOL software provides only the data for the solution and its first time derivative. Therefore, we employ an eighth order finite difference scheme, see [77], in order to approximate accurately the second time derivatives of the solution. Then we proceed with obtaining the reduced basis by carrying out a SVD of the snapshot matrix and thereafter we constructed reduced versions of the displacement and its derivatives. Figure 13 (left) shows the decay of the singular values of the snapshot matrix of the displacement. The cumulative sum of the singular values is shown in Figure 13 (right). We observe that at least the first $50$ singular values are needed for recovering $98$ $\%$ of the energy embedded in the snapshot data.
@@
The operator inference approach used for this example is the time continuous unconstrained version which does not assume full knowledge about the external loads, see Section 3.4. In this approach, we assume only the presence of the input signal (the excitation signal) mentioned earlier. Therefore, the dimension of the control matrix $\widehat{\mathbf{B}}_{M}$ on the right hand side of the reduced system is $r\times 1$.
In this example, we have different scales of the snapshot matrices and the input signal. Therefore, we apply the scaling pre-processing step as explained in Section 3.5. The data matrix of the OpInf problem $\mathbf{\widehat{D}}$ is full rank and therefore one can solve the system of normal equations by inverting the matrix $\mathbf{\widehat{D}}\mathbf{\widehat{D}}^{T}$ to determine the unknown operators. The solution of the linear systems is performed using Matlab 2023b. We infer the reduced operators for different sizes of the reduced model in order to study the effect of the reduced dimension on the quality of the approximation. In particular, we simulate the OpInf problem for $r=10,20,\ldots,100$. We obtained the reduced operators of $\widehat{\mathbf{K}}_{M}$ and $\widehat{\mathbf{B}}_{M}$ for each setting. These reduced operators are then used to integrate the second order mechanical system using the Newmark method. We would like to remark that the results of the reduction given by any set of the learned operators are affected by the choice of the Newmark parameters, which are called $\beta$ and $\gamma$. We refer the reader to [20] in section $5.3$, where two different types of the Newmark method are introduced, namely, the constant average acceleration method ($\beta=\frac{1}{4},\,\gamma=\frac{1}{2}$) and the linear acceleration method ($\beta=\frac{1}{6},\,\gamma=\frac{1}{2}$). We will show the results of the choice which has shown more accuracy which is the linear acceleration method, however, we will also show a comparison between the OpInf-ROM results for both choices for comparison sake.
@@
The solution obtained by the OpInf based ROM and the reference solution of COMSOL are plotted at the sensor location in Figure 14. The plots indicate that the OpInf approach allows for qualitatively accurate approximations of the reference solution for $r=50$ and it can be seen that the reduced approximation is improved when considering $r=80$. A more quantitative assessment of the approach is provided by computing the relative error at the sensor $\epsilon_{s}$ as defined in (66). To analyze the convergence of the OpInf reduced approximation with respect to the reduced dimension, we compute the values of $\epsilon_{s}$ for increasing numbers of reduced basis functions. The values of the sensor error are plotted versus the reduced dimension $r$ in Figure 15 (left) showing that the error reaches smaller values when increasing the reduced dimension $r$.
@@
To further evaluate the quality of the reduced approximation, we also examine the time-dependent relative error over the spatial domain $\epsilon_{u}(t)$, defined in (67). The error $\epsilon_{u}(t)$ for $r=80$ is presented in Figure 15 (right) for two different settings of the Newmark time integrator. The two settings which are described above provide different trends of the error as it can be seen from the plot. This shows that OpInf-ROM accuracy could vary significantly on the choice of the time discretization scheme. The spatial error (of the linear acceleration case) $\epsilon_{u}(t)$ remains under the threshold of 22 $\%$ for the whole part of the simulation which shows that the OpInf-ROM is able to contain the error growth.
@@
Figure 14: Reduced order model solution obtained via unconstrained OpInf for $r=50$ (left) and $r=80$ (right) versus the reference solution of COMSOL at the sensor for the mechanical model of the damaged FML.
@@ 4.3 Numerical results for the nonlinear hyperelastic model
The third numerical example investigates a nonlinear elastic model of an aluminium plate. The model consists of a two-dimensional, intact rectangular aluminium plate with dimensions $0.3$ $\mathrm{m}$ (length) × $0.004$ $\mathrm{m}$ (thickness). The material parameters of aluminium are $\lambda=5.1\times 10^{10}\,$\mathrm{P}\mathrm{a}$\$ and $\mu=2.6\times 10^{10}\,$\mathrm{P}\mathrm{a}$\$. A sketch of the model is shown in Figure 16. We note that this model has been investigated in our article [29], where we considered antisymmetric lamb wave A0. In contrast, here we consider a symmetric lamb wave mode S0. Actuators are positioned at the top and bottom left corners of the plate, applying a five-cycle Hanning window sinusoidal burst with a central frequency of $100$ kHz and a magnitude of excitation which is equal to $10^{-9}$. The material is modeled as hyperelastic, with nonlinearity captured using the Neo-Hookean model [73]. The nonlinear effect of the Neo-Hookean material model can be seen e.g. in [72, Figure 5.1]. The model employs displacement-based excitation, implemented as nonhomogeneous Dirichlet boundary conditions for the vertical displacement at the two left ends of the plate. A homogeneous Dirichlet boundary condition is set for the mid-point on the right side of the plate. The simulation is run for $2.5\times 10^{-4}$ $\mathrm{s}$, with excitation active during the initial $5\times 10^{-5}$ $\mathrm{s}$. The simulation of the GUW is obtained by COMSOL-Multiphysics software©.
Figure 16: Aluminium model setup for the GUW simulation. The red arrows are used to indicate the sign of the excitation signal at each of the two nodes. A symmetric excitation is used for this numerical example through the displacement.
@@
Standard DMD with a reduced dimension of $r=80$ is applied to the same snapshot set, with initial amplitudes computed using equations (61) and (62). However, the method fails in both cases, as demonstrated in Figure 19 (left). This figure compares the out-of-plane displacement measured by the sensor for the reference COMSOL solution and the DMD approximations; standard DMD does not provide even qualitatively accurate results. This does not change as we increase the reduced dimension $r$. In contrast, Figure 19 (right) shows a similar comparison using mrDMD, where the alignment between the signals is significantly better. The relative error $\epsilon_{s}$ for the mrDMD approximation is $1.023$ $\%$, which is within an acceptable range.
@@
In the final numerical test presented in this section, we consider a damage scenario in the aluminium plate, where the damage is modelled as a reduction of the Young’s modulus of elasticity $E$ within a rectangular subdomain. The Young’s modulus of the undamaged aluminum is $E=72\times 10^{9}$ $\mathrm{P}\mathrm{a}$. In the damaged region, the reduced modulus $E_{d}$ is set to $10^{4}$ $\mathrm{P}\mathrm{a}$. The damaged zone is defined by the subdomain $0.06,0.07]\times[0.00124,0.00136]$. The model is then simulated, and displacement snapshots are extracted in order to apply the mrDMD method. In this example, we employ $8$ levels of mrDMD. The results of the full order model (FOM) and the mrDMD reconstruction at the sensor location are shown in (20). The figure demonstrates that mrDMD is able to qualitatively reconstruct the FOM solution. Quantitatively, the sensor error $\epsilon_{s}$ is equal to $4.388$ $\%$.
Figure 20: Comparison between the FOM and mrDMD predicted displacement at the sensor for the nonlinear aluminium model with damage.
@@ 5 Outlook
In this work, we apply both intrusive and non-intrusive model reduction approaches to the wave equation posed on a domain containing damage. Subsequently, we apply non-intrusive reduction techniques to discretized second order mechanical systems obtained from guided ultrasonic wave (GUW) simulations. The test cases considered include models with damage and / or nonlinear behavior.
Damage identification and characterization — determining both the location and severity of defects — are central challenges in the field of structural health monitoring (SHM). This motivation underlies our current research to develop a reduced order model (ROM) capable of generalizing across different damage configurations. To achieve this, we parameterize the damage and address the associated inverse problem, which involves estimating unknown damage parameters from observed system outputs. In this case, the damage parameters influence the full order stiffness matrix and, consequently, the displacement field. This dependency is carried over to the projected reduced operator of the stiffness matrix, i.e. $\mathbf{K}_{r}=\mathbf{K}_{r}(\bm{\theta})$. Since the reduced stiffness matrix is not parameter separable in this case, for each new parameter $\bm{\theta}$ the construction of the reduced stiffness matrix $\mathbf{K}_{r}(\bm{\theta})$ requires the assembly of the full order stiffness matrix $\mathbf{K}(\bm{\theta})$. Thus, the reduced model remains coupled to the high dimension $N$. This bottleneck for multi-query scenarios is well-known such that hyper reduction techniques are developed like matrix gappy POD [17] or matrix DEIM [82]. However, applying such methods requires that the full order solver or computational code allows direct computation of specific entries of the stiffness matrix without assembling the entire matrix. This is not feasible in our case, since we employ COMSOL Multiphysics as the simulator for the GUW problem, and direct access to the full order matrices or the time discretization scheme is restricted. Consequently, we extend the existing OpInf-based ROM framework to handle parameterized systems. The proposed extension must be designed to accurately capture the dependence of the system output s - in this case, the displacement field — on the damage-related input parameters.